OpenStudy (hannahc234):

HannahC234 Please help me! Will fan and medal! How many real number solutions does the equation have? 0=5x^2+2x-12 A. One solution B. Two solutions C. infinitely many solutions D. No solutions

OpenStudy (08surya):

If the discriminant is positive--if b 2 -4ac > 0 --then the quadratic equation has two solutions. If the discriminant is zero--if b 2 - 4ac = 0 --then the quadratic equation has one solution. If the discriminant is negative--if b 2 -4ac < 0 --then the quadratic equation has no solutions. Example: How many solutions does the quadratic equation 2x 2 + 5x + 2 have? a = 2 , b = 5 , and c = 2 . b 2 -4ac = 52 -4(2)(2) = 25 - 16 = 9 > 0 . Thus, the quadratic equation has 2 solutions.

OpenStudy (08surya):

please let me know if you are still not able to solve

OpenStudy (ijlal):

As you can see in the graph the graph intersects the x-axis at 2 points thus it has 2 solutions